Suppose that stars were born at random times over the last 1010 years. The rate of star formation is simply the number of stars divided by 1010 years. The fraction of stars with detected extrasolar planets is at least 20 %. The rate of star formation can be multiplied by this fraction to find the rate planet formation. How often (in years) does a planetary system form in our galaxy? Assume the Milky Way contains 9 × 1011 stars
Answer #1

Number of stars Born Per year is

=(9*1011)/(1010) =90 stars/years

The fraction of stars with detected extrasolar planets is

=(20/100)*(90)=18

Therefore time taken to form a planetary system form in our galaxy is

=1/18 =0.05556 years